# Listening to the Universe with Next Generation Ground-Based Gravitational-Wave Detectors

### v2

@@ VII.2 Limitations of the study
The main caveat of this study is the use of the Fisher information formalism to provide measurement quality estimates. The formalism is well-known and tested, but can only provide estimates for Gaussian posteriors which is likely not the best assumption for the noise of these detectors; especially for signals at visibility threshold. Further, the reliance on numerical derivatives for LAL waveforms and the numerical inversion of the Fisher matrix are sources for numerical uncertainties affecting the quoted results. Further, we did not examine the quality of spin measurements since the addition of such parameters in the Fisher analysis lead to a high rate of ill-conditioned Fisher matrices for which the numerical inversion is not to be trusted.
Lastly, the chosen population distributions—chosen to be consistent with the LIGO and Virgo observations both in mass and redshift distributions—do not capture unexpected sources such as large merger populations of population III star remnants or primordial BHs at large redshifts beyond $z=10$. Besides, we did not include the effects of precession in the examined waveforms and populations.
@@ Acknowledgements.
We thank the members of the Cosmic Explorer project team, in particular Stefan Ballmer, Duncan Brown, Matthew Evans, Anuradha Gupta, Evan Hall, Kevin Kuns, Philippe Landry, Geoffrey Lovelace, Jocelyn Read, David Shoemaker, Josh Smith, Rory Smith, Varun Srivastava, Salvatore Vitale, for many helpful conversations. We thank Ryan Magee in aiding us to verify and correct the sky area estimation of gwbench. We further thank Marica Branchesi, Samuele Ronchini, Michele Maggiore, Francesco Iacovelli, Michele Mancarella, and Stefano Foffa for confirming that the network efficiency and sky resolution of a network containing ET reported in this study is in good agreement with their own findings. SB acknowledges support from NSF grant PHY-1836779 and the Deutsche Forschungsgemeinschaft (DFG), project MEMI number BE6301/2-1. BSS was supported in part by NSF grant PHY-1836779, PHY-2012083 and AST-2006384. We thank all front line workers combating the COVID-19 pandemic without whose support this work would not have been possible.
@@ Appendix D Visibility and measurement quality: histograms
[…]
@@
Figure 11: Histograms of the 10 years BNS injection sample used to generate Fig. 3 for SNR $\rho$, 90%-credible sky area $\Omega_{90}$, fractional errors on chripmass $\Delta\mathcal{M}/\mathcal{M}$ and luminosity distance $\Delta D_{L}/D_{L}$, and absolute errors on symmetric mass ratio $\Delta\eta$ and the inclination angle $\Delta\iota$ observed in the six studied A+, Voyager, and NG networks. The non-SNR panels are obtained for events with SNR $\rho\geq 10$, indicated by the non-shaded region in the SNR panel.
Figure 12: Histograms of the 10 years BBH injection sample used to generate Fig. 4 for SNR $\rho$, 90%-credible sky area $\Omega_{90}$, fractional errors on chripmass $\Delta\mathcal{M}/\mathcal{M}$ and luminosity distance $\Delta D_{L}/D_{L}$, and absolute errors on symmetric mass ratio $\Delta\eta$ and the inclination angle $\Delta\iota$ observed in the six studied A+, Voyager, and NG networks. The non-SNR panels are obtained for events with SNR $\rho\geq 10$, indicated by the non-shaded region in the SNR panel.
@@ Appendix E Cosmic Explorer: Influence of Proposed Detector Configurations
In developing the science case for a NG GW detector proposal, the Cosmic Explorer Project is investigating various configurations of CE detectors, e.g. with varying detector arm lengths and hence detectors sensitivity curves, see Fig. 13. Given the early state of the CE detector proposal, we included, for completeness, four networks containing one or two CE detectors with either $20\,{\rm km}$ or $40\,{\rm km}$ arms to demonstrate the potential of NG networks without an ET detector: VKI+C..20, VKI+C..40, VKI+C..20-40, and VKI+C..40-40. The VKI+C..40 network is the same as VKI+C and is included as a reference from the main body of this paper. Figs. 14 to 22 reiterate the plots presented throughout this paper for these four CE networks instead of the six studied A+, Voyager, and NG networks.
@@
Figure 14: Detection efficiencies $\epsilon$ and detection rates $D_{R}$ of the four studied CE networks are plotted as functions of redshift $z$. The circles (squares) denote the values for events with SNR $\rho\geq 10$ ($\rho\geq 100$). The thick, black lines in the rate panels are the cosmic BNS and BBH merger rates, see Sec. III. The fit lines in the efficiency panels are sigmoid fits with $f_{\rm sigmoid}=\Big{(}\frac{1+b}{1+b\,\mathrm{e}^{ax}}\Big{)}^{c}$.
@@
Since early warning detections heavily depend on a network’s capability to detect GWs at an early stage in the binary evolution, the double-CE networks should observe EW BNS events more often than HLKI+Edue to the higher sensitivity of CE, but much less frequently than either KI+EC or ECSwhich combine the strengths of the ET and CE detectors.
@@
Figure 15: Cumulative histograms for the SNR $\rho$, 90%-credible sky area $\Omega_{90}$, fractional errors on chripmass $\Delta\mathcal{M}/\mathcal{M}$ and luminosity distance $\Delta D_{L}/D_{L}$, and absolute errors on symmetric mass ratio $\Delta\eta$ and inclination angle $\Delta\iota$ for BNS mergers observed in the four studied CE networks. The non-SNR panels are obtained for events with SNR $\rho\geq 10$, indicated by the non-shaded region in the SNR panel. The SNR panel is flipped to highlight the behavior for large values.
Figure 16: Histograms of the 10 years BNS injection sample used to generate Fig. 15 for SNR $\rho$, 90%-credible sky area $\Omega_{90}$, fractional errors on chripmass $\Delta\mathcal{M}/\mathcal{M}$ and luminosity distance $\Delta D_{L}/D_{L}$, and absolute errors on symmetric mass ratio $\Delta\eta$ and the inclination angle $\Delta\iota$ observed in the four studied CE networks. The non-SNR panels are obtained for events with SNR $\rho\geq 10$, indicated by the non-shaded region in the SNR panel.
Figure 17: Cumulative histograms for the SNR $\rho$, 90%-credible sky area $\Omega_{90}$, fractional errors on chripmass $\Delta\mathcal{M}/\mathcal{M}$ and luminosity distance $\Delta D_{L}/D_{L}$, and absolute errors on symmetric mass ratio $\Delta\eta$ and inclination angle $\Delta\iota$ for BBH mergers observed in the four studied CE networks. The non-SNR panels are obtained for events with SNR $\rho\geq 10$, indicated by the non-shaded region in the SNR panel. The SNR panel is flipped to highlight the behavior for large values.
Figure 18: Histograms of the 10 years BBH injection sample used to generate Fig. 17 for SNR $\rho$, 90%-credible sky area $\Omega_{90}$, fractional errors on chripmass $\Delta\mathcal{M}/\mathcal{M}$ and luminosity distance $\Delta D_{L}/D_{L}$, and absolute errors on symmetric mass ratio $\Delta\eta$ and the inclination angle $\Delta\iota$ observed in the four studied CE networks. The non-SNR panels are obtained for events with SNR $\rho\geq 10$, indicated by the non-shaded region in the SNR panel.
@@
Bottom: Cumulative histograms for the SNR $\rho$, 90%-credible sky area $\Omega_{90}$, fractional luminosity distance errors $\Delta D_{L}/D_{L}$, and absolute errors on the inclination angle $\Delta\iota$ for BNS mergers observed in the four studied CE networks for redshifts $z\leq 0.5$. The non-SNR panels are obtained for events with SNR $\rho\geq 10$, indicated by the non-shaded region in the SNR panel. The SNR panel is flipped to highlight the behavior for large values.
@@
Bottom: Cumulative histograms for the SNR $\rho$, 90%-credible sky area $\Omega_{90}$, fractional luminosity distance errors $\Delta D_{L}/D_{L}$, and absolute errors on the inclination angle $\Delta\iota$ for BBH mergers observed in the four studied CE networks for redshifts $z\leq 0.5$. The non-SNR panels are obtained for events with SNR $\rho\geq 10$, indicated by the non-shaded region in the SNR panel. The SNR panel is flipped to highlight the behavior for large values.
Figure 22: Cumulative histograms for the SNR $\rho$ and 90%-credible sky area $\Omega_{90}$ for BNS mergers observed in the four studied CE networks for redshifts $z\leq 0.5$ 2, 5, and 10 minutes before merger. The $\Omega_{90}$ panel is obtained for events with SNR $\rho\geq 10$, indicated by the non-shaded region in the SNR panel. The SNR panel is flipped to highlight the behavior for large values.
@@
[…]
Metric   $\Omega_{90}$ $({\rm deg^{2}})$   $\Delta D_{L}/D_{L}$
Quality   $\leq 1$ $\leq 0.1$ $\leq 0.01$  $\leq 0.1$ $\leq 0.01$
@@
Metric   $\Omega_{90}$ $({\rm deg^{2}})$   $\Delta D_{L}/D_{L}$
Quality   $\leq 1$ $\leq 0.1$ $\leq 0.01$  $\leq 0.1$ $\leq 0.01$
@@
[…]
VKI+C..20 1 0 0 0 0 0 0 0 0 0 0 0
VKI+C..40 2 0 0 0 0 0 0 0 0 0 0 0
VKI+C..20-40 170 21 9 0 35 5 2 0 1 0 0 0
VKI+C..40-40 360 53 20 1 97 13 6 0 4 1 0 0
Figure 23: Histograms of the SNR $\rho_{{\rm post}\textnormal{-}{\rm merger}}$ in the post-merger signal of the BBH population in the four studied CE networks for redshifts $z\leq 0.5$ and a observation time of 10 years.

### v3

@@ VII.2 Limitations of the study
The main caveat of this study is the use of the Fisher information formalism to provide measurement quality estimates. The formalism is well-known and tested, see Vallisneri:2007ev for a review of its shortcomings and Wang:2022kia ; Dupletsa:2024gfl for two approaches to improve the reliability of the Fisher-based error estimates based on Derivative Approximation for Likelihoods and the incorporation of priors for the estimates, respectively. Yet, the formalism only provides estimates for Gaussian posteriors which is likely not the best assumption for the noise of these detectors; especially for signals at visibility threshold. Further, the reliance on numerical derivatives for LAL waveforms and the numerical inversion of the Fisher matrix are sources for numerical uncertainties affecting the quoted results. Further, we did not examine the quality of spin measurements since the addition of such parameters in the Fisher analysis lead to a high rate of ill-conditioned Fisher matrices for which the numerical inversion is not to be trusted.
Lastly, the chosen population distributions—chosen to be consistent with the LIGO and Virgo observations both in mass and redshift distributions—do not capture unexpected sources such as large merger populations of population III star remnants or primordial BHs at large redshifts beyond $z=10$. Besides, the chosen populations only include injections for which the component spins and the binary’s orbital angular momentum are aligned and the chosen waveforms are non-precessing; thus the effects of spin-precession are not accounted for in the forecasts made in this study.
@@ Acknowledgements.
We thank the members of the Cosmic Explorer project team, in particular Stefan Ballmer, Duncan Brown, Matthew Evans, Anuradha Gupta, Evan Hall, Kevin Kuns, Philippe Landry, Geoffrey Lovelace, Jocelyn Read, David Shoemaker, Josh Smith, Rory Smith, Varun Srivastava, Salvatore Vitale, for many helpful conversations. We thank Ryan Magee in aiding us to verify and correct the sky area estimation of gwbench. We further thank Marica Branchesi, Samuele Ronchini, Michele Maggiore, Francesco Iacovelli, Michele Mancarella, and Stefano Foffa for confirming that the network efficiency and sky resolution of a network containing ET reported in this study is in good agreement with their own findings. S. B. acknowledges support from NSF Grant No. PHY-1836779, the Deutsche Forschungsgemeinschaft (DFG), Project MEMI No. BE6301/2-1, ERC Starting Grant No. 945155–GWmining, Cariplo Foundation Grant No. 2021-0555, MUR PRIN Grant No. 2022-Z9X4XS, MUR Grant “Progetto Dipartimenti di Eccellenza 2023-2027” (BiCoQ), and the ICSC National Research Centre funded by NextGenerationEU. B. S. S. was supported in part by NSF Grants No. PHY-1836779, No. PHY-2012083, No. AST-2006384, No. PHY-2207638, No. AST- 2307147, No. PHY-2308886, and No. PHYS-2309064.
@@ Appendix D Visibility and measurement quality: histograms
While events at the tail end of the distribution are rate, they would be very loud and their parameters will be measured with great precision. For example, only NG observatories have the potential to observe significant number of BNS (BBH) mergers with SNRs larger than 100 (1000) and localize sources to better than 10 arc min2. We examine the science potential of these tail ends events in Sec. VI.
@@
Figure 11: Histograms of the 10 years BNS injection sample used to generate Fig. 3 for SNR $\rho$, 90%-credible sky area $\Omega_{90}$, fractional errors on chripmass $\Delta\mathcal{M}/\mathcal{M}$ and luminosity distance $\Delta D_{L}/D_{L}$, and absolute errors on symmetric mass ratio $\Delta\eta$ and the inclination angle $\Delta\iota$ observed in the six studied A+, Voyager, and NG networks. The histograms were generated from $\sim 4.7\times 10^{6}$ injections sampled according to Sec. IV.1. The non-SNR panels are obtained for events with SNR $\rho\geq 10$, indicated by the non-shaded region in the SNR panel.
Figure 12: Histograms of the 10 years BBH injection sample used to generate Fig. 4 for SNR $\rho$, 90%-credible sky area $\Omega_{90}$, fractional errors on chripmass $\Delta\mathcal{M}/\mathcal{M}$ and luminosity distance $\Delta D_{L}/D_{L}$, and absolute errors on symmetric mass ratio $\Delta\eta$ and the inclination angle $\Delta\iota$ observed in the six studied A+, Voyager, and NG networks. The histograms were generated from $\sim 1.2\times 10^{6}$ injections sampled according to Sec. IV.1. The non-SNR panels are obtained for events with SNR $\rho\geq 10$, indicated by the non-shaded region in the SNR panel.
@@ Appendix E Cosmic Explorer: Influence of Proposed Detector Configurations
In developing the science case for a NG GW detector proposal, the Cosmic Explorer Project is investigating various configurations of CE detectors, e.g. with varying detector arm lengths and hence detectors sensitivity curves, see Fig. 13. Given the early state of the CE detector proposal, we included, for completeness, four networks containing one or two CE detectors with either $20\,{\rm km}$ or $40\,{\rm km}$ arms to demonstrate the potential of NG networks without an ET detector: VKI+C..20, VKI+C..40, VKI+C..20-40, and VKI+C..40-40. The VKI+C..40 network is the same as VKI+C and is included as a reference from the main body of this paper. Figs. 14 to 23 reiterate the plots presented throughout this paper for these four CE networks instead of the six studied A+, Voyager, and NG networks.
@@
Figure 14: Detection efficiencies $\epsilon$ and detection rates $D_{R}$, see Eqs. (3) and (9), respectively, of the four studied CE networks are plotted as functions of redshift $z$. The efficiency specifies the value at the indicated redshift $z$ while the detection rate corresponds to the integrated rate up to that redshift. The circles (squares) denote the values for events with SNR $\rho\geq 10$ ($\rho\geq 100$). The thick, black lines in the rate panels are the cosmic BNS and BBH merger rates, see Sec. III. The fit lines in the efficiency panels are sigmoid fits with $f_{\rm sigmoid}=\Big{(}\frac{1+b}{1+b\,\mathrm{e}^{ax}}\Big{)}^{c}$.
@@
Since early warning detections heavily depend on a network’s capability to detect GWs at an early stage in the binary evolution, the double-CE networks should observe EW BNS events more often than HLKI+Edue to the higher sensitivity of CE, but much less frequently than either KI+EC or ECSwhich combine the strengths of the ET and CE detectors (see Table 15 in comparison to Table 8).
@@
Figure 15: Cumulative histograms for the SNR $\rho$, 90%-credible sky area $\Omega_{90}$, fractional errors on chripmass $\Delta\mathcal{M}/\mathcal{M}$ and luminosity distance $\Delta D_{L}/D_{L}$, and absolute errors on symmetric mass ratio $\Delta\eta$ and inclination angle $\Delta\iota$ for BNS mergers observed in the four studied CE networks. The histograms were generated from $\sim 4.7\times 10^{6}$ injections sampled according to Sec. IV.1. The non-SNR panels are obtained for events with SNR $\rho\geq 10$, indicated by the non-shaded region in the SNR panel. The SNR panel is flipped to highlight the behavior for large values.
Figure 16: Histograms of the 10 years BNS injection sample used to generate Fig. 15 for SNR $\rho$, 90%-credible sky area $\Omega_{90}$, fractional errors on chripmass $\Delta\mathcal{M}/\mathcal{M}$ and luminosity distance $\Delta D_{L}/D_{L}$, and absolute errors on symmetric mass ratio $\Delta\eta$ and the inclination angle $\Delta\iota$ observed in the four studied CE networks. The histograms were generated from $\sim 4.7\times 10^{6}$ injections sampled according to Sec. IV.1. The non-SNR panels are obtained for events with SNR $\rho\geq 10$, indicated by the non-shaded region in the SNR panel.
Figure 17: Cumulative histograms for the SNR $\rho$, 90%-credible sky area $\Omega_{90}$, fractional errors on chripmass $\Delta\mathcal{M}/\mathcal{M}$ and luminosity distance $\Delta D_{L}/D_{L}$, and absolute errors on symmetric mass ratio $\Delta\eta$ and inclination angle $\Delta\iota$ for BBH mergers observed in the four studied CE networks. The histograms were generated from $\sim 1.2\times 10^{6}$ injections sampled according to Sec. IV.1. The non-SNR panels are obtained for events with SNR $\rho\geq 10$, indicated by the non-shaded region in the SNR panel. The SNR panel is flipped to highlight the behavior for large values.
Figure 18: Histograms of the 10 years BBH injection sample used to generate Fig. 17 for SNR $\rho$, 90%-credible sky area $\Omega_{90}$, fractional errors on chripmass $\Delta\mathcal{M}/\mathcal{M}$ and luminosity distance $\Delta D_{L}/D_{L}$, and absolute errors on symmetric mass ratio $\Delta\eta$ and the inclination angle $\Delta\iota$ observed in the four studied CE networks. The histograms were generated from $\sim 1.2\times 10^{6}$ injections sampled according to Sec. IV.1. The non-SNR panels are obtained for events with SNR $\rho\geq 10$, indicated by the non-shaded region in the SNR panel.
@@
Bottom: Cumulative histograms for the SNR $\rho$, 90%-credible sky area $\Omega_{90}$, fractional luminosity distance errors $\Delta D_{L}/D_{L}$, and absolute errors on the inclination angle $\Delta\iota$ for BNS mergers observed in the four studied CE networks for redshifts $z\leq 0.5$. The histograms were generated from $\sim 1.2\times 10^{5}$ injections sampled according to Sec. IV.1. The non-SNR panels are obtained for events with SNR $\rho\geq 10$, indicated by the non-shaded region in the SNR panel. The SNR panel is flipped to highlight the behavior for large values.
@@
Bottom: Cumulative histograms for the SNR $\rho$, 90%-credible sky area $\Omega_{90}$, fractional luminosity distance errors $\Delta D_{L}/D_{L}$, and absolute errors on the inclination angle $\Delta\iota$ for BBH mergers observed in the four studied CE networks for redshifts $z\leq 0.5$. The histograms were generated from $\sim 1.2\times 10^{4}$ injections sampled according to Sec. IV.1. The non-SNR panels are obtained for events with SNR $\rho\geq 10$, indicated by the non-shaded region in the SNR panel. The SNR panel is flipped to highlight the behavior for large values.
Figure 22: Cumulative histograms for the SNR $\rho$ and 90%-credible sky area $\Omega_{90}$ for BNS mergers observed in the four studied CE networks for redshifts $z\leq 0.5$ 2, 5, and 10 minutes before merger. The histograms were generated from $\sim 1.2\times 10^{5}$ injections sampled according to Sec. IV.1. The $\Omega_{90}$ panel is obtained for events with SNR $\rho\geq 10$, indicated by the non-shaded region in the SNR panel. The SNR panel is flipped to highlight the behavior for large values.
@@
Metric AAA     $\Omega_{90}$ $({\rm deg^{2}})$ A     $\Delta D_{L}/D_{L}$  A
Quality     $\leq 1$ $\leq 0.1$ $\leq 0.01$A    $\leq 0.1$ $\leq 0.01$
@@
[…]
Metric     $\Omega_{90}$ $({\rm deg^{2}})$ A     $\Delta D_{L}/D_{L}$  A
Quality     $\leq 1$ $\leq 0.1$ $\leq 0.01$A    $\leq 0.1$ $\leq 0.01$
@@
$\Omega_{90}$ (deg2)ABC $\leq 100$ $\leq 20$ $\leq 10$ $\leq 1$ A $\leq 100$ $\leq 20$ $\leq 10$ $\leq 1$ A $\leq 100$ $\leq 20$ $\leq 10$ $\leq 1$ A
VKI+C..20 1 0 0 0 ABC 0 0 0 0 ABC 0 0 0 0
VKI+C..40 2 0 0 0 ABC 0 0 0 0 ABC 0 0 0 0
VKI+C..20-40 170 21 9 0 ABC 35 5 2 0 ABC 1 0 0 0
VKI+C..40-40 360 53 20 1 ABC 97 13 6 0 ABC 4 1 0 0
Figure 23: Histograms of the SNR $\rho_{{\rm post}\textnormal{-}{\rm merger}}$ in the post-merger signal of the BBH population in the four studied CE networks for redshifts $z\leq 0.5$ and a observation time of 10 years. The histograms were generated from $\sim 1.2\times 10^{4}$ injections sampled according to Sec. IV.1.